II Problems (4 points each) 7. Assume you are the consultant of an online movie-rental company. Suppose th year cash flow assoc stay with the company for 3 more years generating an annual cash company has a 30% chance of retaining a new subscriber. Given a 12% discount rate, what is the expected value of a new subscriber? (Assume all cash flows iated with a new subscriber is -$20. If retained, the subscriber will flow of $100. The occur at the end of the year.)
Answer #1

present value of retained amount = 100/1.12 + 100/1.12^2 + 100/1.12^3 = 240.18

expected value of new subscriber

= probability of retaining * present value of retained amount + probability of not retained * value of not retaining

= 0.3 * 240.18 + 0.7 * -20

= 58.05